$SPAD/src/input c06fqf.input

The Axiom Team December 3, 2016

Abstract

Contents

2

1 License

— license —

--Copyright The Numerical Algorithms Group Limited 1994.

———-

— * —

\getchunk{license}
)clear all
showArrayValues true
showScalarValues true
x:Matrix SF:=
 [[0.3854 ,0.6772 ,0.1138 ,0.6751 ,0.6362 ,0.1424],_
  [0.5417 ,0.2983 ,0.1181 ,0.7255 ,0.8638 ,0.8723],_
  [0.9172 ,0.0644 ,0.6037 ,0.6430 ,0.0428 ,0.4815 ]]
m:=nrows x
n:=ncols x
trig:Matrix SF:=new(1,2*n,0.0@SF)
init:="i"
result:=c06fqf(m,n,init,x,trig,-1)

———-

References^{[1] nothing}

4